# Automated Counting of Spanning Trees for Several Infinite families of Graphs

### v2

@@ 5.2 Generating functions for the Total Number of Leaves
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{G}_{2}$ is
@@
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{G}_{3}$ is
@@
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{G}_{4}$ is
@@ 5.3 B-Z Constants
See Section 4 to read about our method for computing these constants. Recall that $BZ(\mathcal{G})$ denotes the B-Z constant of a graph family $\mathcal{G}$.
@@
The B-Z constant for $\mathcal{G}_{2}$ is
$-6+\frac{14\sqrt{5}}{5}.$
@@
Let $\alpha$ be the smallest real root of $z^{8}-4z^{7}-17z^{6}+8z^{5}+49z^{4}+8z^{3}-17z^{2}-4z+1$. Then, the B-Z constant for $\mathcal{G}_{4}$ is
@@
$BZ(\mathcal{G}_{3})<\frac{117451}{355635}<BZ(\mathcal{G}_{4})<\frac{117452}{355635}$
@@ 6.1 Generating functions for the Number of Spanning Trees
In this section, we include our results for the Number of Spanning Trees in the class $\mathcal{H}_{r}$, with $2\leq r\leq 6$.
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{H}_{2}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{H}_{3}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{H}_{4}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{H}_{5}$ is
@@ 6.2 Generating functions for the Total Number of Leaves
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{H}_{2}$ is
@@
The generating function for the total number of leaves in $\mathcal{H}_{3}$ is
@@
The generating function for the total number of leaves in $\mathcal{H}_{4}$ is
$\frac{C_{4}}{(t^{6}-3t^{5}+6t^{4}-10t^{3}+6t^{2}-3t+1)^{2}(t^{8}-4t^{7}-17t^{6}+8t^{5}+49t^{4}+8t^{3}-17t^{2}-4t+1)^{2}}$
where $C_{4}$ is a degree $29$ polynomial. Note that the degree of the numerator is $28$
@@ 6.3 B-Z Constants
See Section 4 to read about our method for computing these constants. Compare the results in this section to Section 5.3.
@@
The B-Z constant for $\mathcal{H}_{2}$ is
$-6+\frac{14\sqrt{5}}{5}.$
@@
$BZ(\mathcal{H}_{4})=BZ(\mathcal{G}_{4})$.
We also verified that $BZ(\mathcal{H}_{5})=BZ(\mathcal{G}_{5})$. See [1].
@@ 7 B-Z Constants for Grid and Torus Graphs
In previous sections, we discussed powers of cycles and powers of paths. The members of these two graph families were related by a subgraph relation and we found their B-Z constants to be the same up to $r=5$. Since a power of a cycle is vertex-transitive, computing the total number of leaves was faster (Corollary 3.0.2). In this section, we discuss another pair of graph families – Grid Graphs and Torus Graphs – which are also related by a subgraph relation. Similarly to the previous pair of families, Torus graphs are vertex-transitive. However, the B-Z constants for grid graphs and torus graphs differ.
@@
The B-Z constant for: $4\times n$ torus graphs is $0.2917148\ldots$; $5\times n$ torus graphs is $0.29342497\ldots$; and $6\times n$ torus graphs is $0.294011\ldots$. The B-Z constant for $4\times n$ grid graphs is .$2746417\ldots.$.
We computed B-Z constants for fewer grid families (than torus families) because grid graphs are not vertex-transitive; hence, we were unable to apply Corollary 3.0.2 to speed up computation of the "total number of leaves".
8 Verification of Results by Random Sampling
In our implementation, we wrote a procedure to compute the total number of leaves of a graph (NumLeaves or VtxTransNumLeaves) and another to compute the number of spanning trees (NumSpanTree). The ratio between the outputs of these procedures is the average number of leaves of a graph.
To verify the accuracy of these procedures, we selected a large member from each of our graph families (at least $100$ vertices) and then sampled a large number of spanning trees from that graph (at least $100$ and usually more than $200$). After doing so, we computed the average number of leaves of the graph in the sample and compared it to our procedures’ exact computation. In all cases, we found that the estimate and the true values were similar. For more detailed results on this verification, we direct the reader to [1].
The algorithm we used to sample uniformly random spanning trees is Wilson’s Algorithm, as described in [3].
@@ 9 Accompanying Maple package
NumSpanTree(n,E) given a positive integer n and a set of edges E on the set $\{1,\ldots,n\}$, the procedure returns the number of spanning trees of the corresponding graph $n,E$.
NumSpanTreeSeq(F, ArgList, a,b) given the name of a graph-generating procedure F and a list of arguments ArgList, outputs a list of the number of spanning trees for the graphs F(a, op(ArgList))$\ldots$,F(b, op(ArgList)).
@@
NumLeavesSeq(F, ArgList, a,b) given the name of a graph-generating procedure F and a list of arguments ArgList, outputs a list of the total number of leaves for the graphs F(a, op(ArgList))$\ldots$,F(b, op(ArgList)).
VtxTransNumLeavesSeq(F, ArgList, a,b) similar to NumLeavesSeq. However, assumes that the graphs generated by F are vertex-transitive. Uses Corollary 3.0.2 to compute the output faster.
@@ 10 Conjectures and Further Discussion
    For each $r$ we computed, the denominator of a rational function from Section 6.1 divides the (corresponding) denominator of a rational function from Section 5.1. Similarly for Section 6.2 and Section 5.2.
@@
    In the denominators for the generating functions given in Section 5.1, Section 5.2, Section 6.1, Section 6.2, each polynomial factor has coefficients which are symmetric. Is this also true for all other recursive graph families?
@@
The B-Z constants of $\mathcal{G}_{r}$ and $\mathcal{H}_{r}$ are equal for all $r\geq 1$.
Assuming that the previous conjecture holds, a natural question arises: given a graph family, is there a threshold for how many vertices you can "modify" (by edge deletion) while keeping the same B-Z constant? It turns out that Grid graphs and Torus Graphs have different B-Z constants, which may suggest that we can only modify a finite number of vertices. Originally, we considered the following conjecture. Note that the conditions in the following conjecture are satisfied when $\mathcal{G}=\mathcal{G}_{r}$ and $\mathcal{H}=\mathcal{H}_{r}$ (as defined in Section 2).
@@
(False) Let $\mathcal{G}:=\{G_{n}\}_{n}$ and :=$\mathcal{H}:=\{H_{n}\}_{n}$ be (indexed) families of connected graphs, such that $BZ(\mathcal{G})$ and $BZ(\mathcal{H})$ both exist. If the following hold:
@@
Counterexample. Let $\mathcal{H}$ be the family of graphs constructed in the proof of Proposition 4.0.2 with B-Z constant equal to $\frac{1}{2}$. In other words, $\mathcal{H}$ is the family of graphs resulting from subdividing each edge in a star graph once. From each $H_{n}\in\mathcal{H}$, construct $G_{n}$ by adding an edge between two non-leaf vertices. We see that $\tau(G_{n})=3\cdot\tau(H_{n})$, but $\mathcal{L}(G_{n})=\mathcal{L}(H_{n})$. Hence, $BZ(\mathcal{G})=\frac{1}{3}\cdot BZ(\mathcal{H})=\frac{1}{6}\neq\frac{1}{2}=BZ(\mathcal{H})$.
@@
Let $n\geq 2r+1$. If $G_{n}\in\mathcal{G}_{r}$ and $H_{n}\in\mathcal{H}_{r}$, then $G_{n}\setminus H_{n}$ contains a (copy of the) Half Graph on $2r$ vertices and $n-2r$ isolated vertices.
The Half Graph on $2n$ vertices is a bipartite graph with vertex set $\{u_{1},\ldots,u_{n},v_{1},\ldots,v_{n}\}$ and edges $\{u_{i}v_{j}\ :\ i\leq j\}$. Since $r$ is fixed, the number of differing edges between a Power of a Cycle and a Power of a Path is finite.

### v3

@@ 5.2 Generating functions for the Total Number of Leaves
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{C}_{2}$ is
@@
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{C}_{3}$ is
@@
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{C}_{4}$ is
@@ 5.3 B-Z Constants
See Section˜4 to read about our method for computing these constants. Recall that $BZ(\mathcal{G})$ denotes the B-Z constant of a graph family $\mathcal{G}$.
@@
The B-Z constant for $\mathcal{C}_{2}$ is
$6+\frac{14\sqrt{5}}{5}.$
@@
Let $\alpha$ be the smallest real root of $z^{8}-4z^{7}-17z^{6}+8z^{5}+49z^{4}+8z^{3}-17z^{2}-4z+1$. Then, the B-Z constant for $\mathcal{C}_{4}$ is
@@
$BZ(\mathcal{C}_{3})<\frac{117451}{355635}<BZ(\mathcal{C}_{4})<\frac{117452}{355635}$
@@ 6.1 Generating functions for the Number of Spanning Trees
In this section, we present our results for the Number of Spanning Trees in the class $\mathcal{P}_{k}$, with $2\leq k\leq 6$.
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{P}_{2}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{P}_{3}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{P}_{4}$ is
@@
The generating function $f(t)$ for the number of spanning trees in $\mathcal{P}_{5}$ is
@@ 6.2 Generating functions for the Total Number of Leaves
The generating function for the total number of leaves (across all spanning trees of a member) in $\mathcal{P}_{2}$ is
@@
The generating function for the total number of leaves in $\mathcal{P}_{3}$ is
@@
The generating function for the total number of leaves in $\mathcal{P}_{4}$ is
$\frac{M_{4}}{(t^{6}-3t^{5}+6t^{4}-10t^{3}+6t^{2}-3t+1)^{2}(t^{8}-4t^{7}-17t^{6}+8t^{5}+49t^{4}+8t^{3}-17t^{2}-4t+1)^{2}}$
where $M_{4}$ is a degree $29$ polynomial in $t$. Note that the degree of the denominator is $28.$
@@ 6.3 B-Z Constants
See Section˜4 to read about our method for computing these constants. Compare the results in this section to Section˜5.3.
@@
The B-Z constant for $\mathcal{P}_{2}$ is
$6+\frac{14\sqrt{5}}{5}.$
@@
$BZ(\mathcal{P}_{4})=BZ(\mathcal{C}_{4})$.
We also verified that $BZ(\mathcal{P}_{5})=BZ(\mathcal{C}_{5})$. See [2].
@@ 7 B-Z Constants for Grid and Torus Graphs
In previous sections, we discussed powers of cycles and powers of paths. The members of these two graph families were related by a subgraph relation and we found their B-Z constants to be the same up to $r=5$. Since a power of a cycle is vertex-transitive, computing the total number of leaves was faster (Corollary˜3.0.2). In this section, we discuss another pair of graph families – grid graphs and torus graphs – which are also related by a subgraph relation. Similarly to the previous pair of families, torus graphs are vertex-transitive. However, the B-Z constants for grid graphs and torus graphs differ.
@@
The B-Z constant for: $4\times n$ torus graphs is $0.2917148\dots$; $5\times n$ torus graphs is $0.29342497\dots$; and $6\times n$ torus graphs is $0.294011\dots$. Notably, these B-Z constants appear to approach the B-Z constant for $n\times n$ torus graphs – which is remarked to be $(1-\frac{2}{\pi})\cdot\frac{8}{\pi^{2}}$ by [3, Remark 5.3], who furthermore cite [5, p. 112]. The B-Z constant for $4\times n$ grid graphs is $2746417\dots.$
We computed B-Z constants for fewer grid families (than torus families) because grid graphs are not vertex-transitive; hence, we were unable to apply Corollary˜3.0.2 to speed up computation of the “total number of leaves".
8 Estimating B-Z Constants with Random Sampling
As we have shown in previous sections, with enough computational resources, it is possible to find the exact B-Z constant for several graph families. Even though the exact B-Z constant may be given as a root of a polynomial, one can also obtain precise numerical bounds for it. To perform these exact computations, we wrote a procedure to compute the total number of leaves of a graph (NumLeaves or VtxTransNumLeaves) and another to compute the number of spanning trees (NumSpanTree). The ratio between the outputs of these procedures gave us the average number of leaves of a graph.
However, if one only cares about finding a numerical estimate of the B-Z constant for a graph family – after confirming it exists – there is another approach. Fix a member of the graph family with a large number $n>N$ of vertices. Then, sample a large number $K$ of (uniformly) random spanning trees from the graph $G_{n}$, count the number of leaves in each sampled tree, and use the resulting data to estimate the average number of leaves for the large member $G_{n}$ in the family. The B-Z constant is an asymptotic value; so, for large enough $N$ and $K$, a numerical sample will suffice.
In our Maple package [2], we implemented a procedure (EstAvgNumLeaves) that uses parameters $N,K$, as described above, to estimate the average number of leaves in such a graph $G_{n}$. Dividing the output from EstAvgNumLeaves by the number of vertices ($n$) will result in a numerical estimate for the B-Z constant. To sample uniformly random spanning trees, we implemented Wilson’s algorithm, described in [7].
Using the procedure, we compared numerical estimates to the exact B-Z constants for the graph families in this paper. We used $N\geq 100$ and $K\geq 100$ (and usually $K\geq 200$). As expected, we found that the estimated and the true B-Z constants were similar. The data from these numerical estimates can be found in [2].
@@ 9 Accompanying Maple package
The Help() procedure in the Maple package, available at [2], will provide even more details, as well as examples for how to use these procedures.
@@
NumSpanTree(n,E) given a positive integer n and a set of edges E on the set $\{1,\dots,n\}$, the procedure returns the number of spanning trees of the corresponding graph $n,E$.
[…]
@@
[…]
VtxTransNumLeavesSeq(F, ArgList, a,b) similar to NumLeavesSeq. However, assumes that the graphs generated by F are vertex-transitive. Uses Corollary˜3.0.2 to compute the output faster.
@@ 10 Conjectures and Further Discussion
    For each $r$ we computed, the denominator of a rational function from Section˜6.1 divides the (corresponding) denominator of a rational function from Section˜5.1. Similarly for Section˜6.2 and Section˜5.2.
@@
    In the denominators for the generating functions given in Section˜5.1, Section˜5.2, Section˜6.1, Section˜6.2, each polynomial factor has coefficients which are symmetric. Is this also true for all other recursive graph families?
@@
The B-Z constants of $\mathcal{C}_{k}$ and $\mathcal{P}_{k}$ are equal for all $k\geq 1$.
Assuming that the previous conjecture holds, a natural question arises: given a graph family, is there a threshold for how many vertices you can "modify" (by edge deletion) while keeping the same B-Z constant? It turns out that grid graphs and torus graphs have different B-Z constants, which may suggest that we can only modify a finite number of vertices. Originally, we considered the following conjecture. Note that the conditions in the following conjecture are satisfied when $\mathcal{G}=\mathcal{C}_{k}$ and $\mathcal{H}=\mathcal{P}_{k}$ (as defined in Section˜2).
@@
(False) Let $\mathcal{G}:=\{G_{n}\}_{n}$ and $\mathcal{H}:=\{H_{n}\}_{n}$ be (indexed) families of connected graphs, such that $BZ(\mathcal{G})$ and $BZ(\mathcal{H})$ both exist. If the following hold:
@@
Counterexample. Let $\mathcal{H}$ be the family of graphs constructed in the proof of Proposition˜4.0.2 with B-Z constant equal to $\frac{1}{2}$. In other words, $\mathcal{H}$ is the family of graphs resulting from subdividing each edge in a star graph once. From each $H_{n}\in\mathcal{H}$, construct $G_{n}$ by adding an edge between two non-leaf vertices. We see that $\tau(G_{n})=3\cdot\tau(H_{n})$, but $\mathcal{L}(G_{n})=\mathcal{L}(H_{n})$. Hence, $BZ(\mathcal{G})=\frac{1}{3}\cdot BZ(\mathcal{H})=\frac{1}{6}\neq\frac{1}{2}=BZ(\mathcal{H})$.
@@
Let $n\geq 2r+1$. If $G_{n}\in\mathcal{C}_{k}$ and $H_{n}\in\mathcal{P}_{k}$, then $G_{n}\setminus H_{n}$ contains a (copy of the) half graph on $2r$ vertices and n2rn-2r isolated vertices.
The half graph on $2n$ vertices is a bipartite graph with vertex set $\{u_{1},\dots,u_{n},v_{1},\dots,v_{n}\}$ and edges $\{u_{i}v_{j}\ :\ i\leq j\}$. Since $r$ is fixed, the number of differing edges between a power of a cycle and a power of a path is finite.
Acknowledgments
The authors wish to thank an anonymous reviewer for their detailed and insightful comments that improved the readability of this paper, as well as for providing additional references to the paper (which relate it to pre-existing results and literature).